# Growth of (Cu,C)Ba2Ca2Cu3O9+δ thin films on flexible Hastelloy tapes

### v1

@@ Abstract
The applications of superconducting cable or magnet require that the superconductors are made into wires or tapes. For cuprate superconductors, this is a big challenge because of the strong flux motion induced by high anisotropy, very short coherence length and strong thermal fluctuation, etc. One of the ways is to fabricate superconducting films on flexible metallic tapes with oxide buffer layers. The successful one so far is the REBa2Cu3O7 (REBCO, RE=rare earth elements) films in tape form, as called the coated conductors. While the superconducting transition temperature of REBCO system is limited to about 90 K. Here we report the successful fabrication of another new non-toxic superconducting film, namely (Cu,C)Ba2Ca2Cu3O9+δ on these flexible metallic tapes with LaMnO3 and CeO2 as the top layers. The onset superconducting transition occurs at 112 K and 110 K, and the zero-resistance transition temperatures are about 96 K and 98 K, respectively. The temperature dependent resistivity under magnetic fields in different directions reveal a relatively small anisotropy. Further optimization of the films will improve the zero resistance transition temperature, thus can also improve the characteristic properties for applications. Our results show that the (Cu,C)Ba2Ca2Cu3O9+δ is a promising candidate material for the high power applications in liquid nitrogen temperature region.
@@ 1 Introduction
Since the discovery of superconductivity[1], people immediately realized that superconductors may be applied in many ways, thus enormous efforts have been dedicated to explore practical superconductors that meet the industrial requirements. Superconductors can be categorized into low-temperature superconductors (LTS) and high-temperature superconductors (HTS) according to their critical temperatures. Applicable low-temperature superconductors include NbTi, Nb3Sn and so on. NbTi superconductors have very good ductility and can be directly made into long wires[2]. Nb3Sn, on the other hand, due to its brittleness and hardness, is currently prepared by the bronze process and powder-in-tube (PIT) method[3]. Both superconductors are currently commercialized. The critical temperature of some cuprate high temperature superconductors is higher than the liquid nitrogen boiling temperature (77 K), such as Hg, Tl, Bi and RE-based systems[4, 5, 6, 7]. Due to the existence of toxic elements and the high anisotropy, the application of Hg-based and Tl-based systems is limited. Bi-based system, as a ceramic material, can not be directly made into wires and are mainly manufactured by powder-in-tube method, known as the first generation of high-temperature superconducting tapes. Because of its strong anisotropy, the vortex motion is very strong leading to a rather low irreversibility magnetic field in the liquid nitrogen temperature region, and the critical current density also decreases quickly in a small magnetic field at 77 K[8]. The critical temperature of REBa2Cu3O7-δ (REBCO) is about 90 K, and the anisotropy is relatively small compared with the Bi-system[9], which means that there is a high irreversibility field in the liquid nitrogen temperature region. However, due to the very small coherence length and the small tolerant misorientation angle between neighboring grains for the superconducting current to go through, the PIT technique was not successful to make the long wires for REBCO. In order to eliminate the weak links from the grain boundaries and obtain a better grain alignment to meet the requirements for practical applications, REBCO can be fabricated by epitaxial thin film method such as metal-organic chemical vapour deposition(MOCVD) and pulsed laser deposition (PLD)[10], etc. Thus the REBCO is known as the second generation of high-temperature superconducting tapes. However, the critical temperature of REBCO is about 90 K, lower than other systems with three or four layers of $CuO_{2}$ planes in the superconducting block. There is an another high-temperature superconducting system, (Cu,C)Ba2Can-1CunO2n+3 (n=1,2,3, …) [11, 12, 13, 14, 15]. Here (Cu,C)Ba2Ca2Cu3O9+δ((Cu,C)-1223) and (Cu,C)Ba2Ca3Cu4O11+δ((Cu,C)-1234) have a superconducting transition temperature of about 120 K and a higher irreversibility field than REBCO. But (Cu,C)-1223 and (Cu,C)-1234 bulks can only be synthesized by high-pressure methods, strongly limiting their applications. The growth of (Cu,C)-based superconducting films with high transition temperature is still challenging. The transition temperatures of (Cu,C)-1223 and (Cu,C)-1234 films grown on single-crystal substrates have reached 99.7 K and 96 K  [16, 17, 18], respectively. Similar to REBCO, these materials need to be epitaxially grown on flexible metallic substrate to meet the demands of industrial applications. However, to the best of our knowledge, up to now, there are no reports of (Cu,C)-1223 films grown on flexible metallic substrates. Due to its superior properties, it is important to grow this material on flexible metallic substrate in order to meet the demand for the industrial applications.
@@ 2 Experimental details
The (Cu,C)Ba2Ca2Cu3O9+δ thin films were epitaxially grown on Hastelloy tapes with LaMnO3 and CeO2 top buffer layer by pulsed laser deposition technique with a 248-nm KrF excimer laser, using the nominal composition target of (Cu,C)Ba2Ca3Cu4.6Oy. The Hastelloy tapes with LaMnO3 and CeO2 top buffer layer are obtained from Companies of Suzhou Advanced Materials and Shanghai Superconductor, respectively. Fig. 1(a) and (b) show the architectures of the (Cu,C)-1223 coated conductors on different top layers. During deposition, a mixed gas of O2 and CO2 was introduced into the chamber, and the total pressure was maintained at 15-25 Pa. The deposition temperature for all films was controlled in the region 620-680 $\rm{{}^{\circ}C}$. The laser energy density of 1-2 J/cm2 and repetition rate of 5 Hz were used to ablate the target. After the deposition, the films were cooled down to room temperature in deposition pressure at a rate of $5\rm{{}^{\circ}C}/min$.
Figure 1: (a), (b) Sketch of two-type architectures of the (Cu,C)-1223 films. (c), (d) X-ray diffraction (XRD) patterns of the (Cu,C)Ba2Ca2Cu3O9+δ films on flexible Hastelloy substrates with the top layer of LaMnO3 and CeO2, respectively. (e), (f) AFM images of the (Cu,C)-1223/LaMnO3 film and the (Cu,C)-1223/CeO2 film.
@@ 3 Results and discussion
The x-ray diffraction patterns of the (Cu,C)-1223 films on flexible Hastelloy substrates with LaMnO3 and CeO2 top buffer layer are displayed in Fig. 1(c) and (d). All observed (00l) diffraction peaks can be indexed to (Cu,C)-1223 and the substrates, indicating that the films have been grown with c-axis orientation. Analysis of the XRD data yields a lattice constant $c$ = 14.82 Å for both films. Surface morphology and roughness have been investigated through atomic force microscopy (AFM) as shown in Fig. 1(c) and (f). The average roughness (Ra) of the (Cu,C)-1223/LaMnO3 film is about 1.9 nm, determined over the scan area $5\times 5\mu m^{2}$. For the (Cu,C)-1223/CeO2 film, the average roughness is about 5.0 nm. One possible reason for the obvious different $R_{a}$ is that the growth of the (Cu,C)-1223 films (a=3.86 Å) may be affected by epitaxial tensile strain from LaMnO3 (a=3.94 Å) and epitaxial compressive strain from CeO2 (a=3.82 Å).
Figure 2: (a), (b) Temperature dependent resistivity of the (Cu,C)Ba2Ca2Cu3O9+δ films on flexible metallic substrates with the top layer of CeO2 and LaMnO3. (c), (d) The temperature dependence of magnetic susceptibility under a magnetic field of 1 Oe.
@@
Figure 3: (a), (b) Temperature dependent resistivity of the (Cu,C)-1223/LaMnO3 film under magnetic fields $H\parallel c$ and $H\parallel ab$, respectively. (c) Temperature dependence of the upper critical field $H_{c2}$. The solid lines show the linear fitting results. (d) The resistivity as a function of the effective magnetic field at 90 K. The inset shows that angle dependent resistivity under different magnetic fields.

### v2

# Growth of (Cu,C)Ba2Ca2Cu3O9±δ thin films on flexible Hastelloy tapes

@@ Abstract
The applications of superconducting cable or magnet require that the superconductors are made into wires or tapes. For cuprate superconductors, this is a big challenge because of the strong flux motion induced by high anisotropy, very short coherence length and strong thermal fluctuation, etc. One of the ways is to fabricate superconducting films on flexible metallic tapes with oxide buffer layers. The successful one so far is the REBa2Cu3O7 (REBCO, RE=rare earth elements) films in tape form, as called the coated conductors. While the superconducting transition temperature of REBCO system is limited to about 90 K. Here we report the successful fabrication of another new non-toxic superconducting film, namely (Cu,C)Ba2Ca2Cu3O9±δ on these flexible metallic tapes with LaMnO3 and CeO2 as the top layers. The onset superconducting transition occurs at 112 K and 110 K, and the zero-resistance transition temperatures are about 96 K and 98 K, respectively. The temperature dependent resistivity under magnetic fields in different directions reveal a relatively small anisotropy. Further optimization of the films will improve the zero resistance transition temperature, thus can also improve the characteristic properties for applications. Our results show that the (Cu,C)Ba2Ca2Cu3O9±δ is a promising candidate material for the high power applications in liquid nitrogen temperature region.
@@ 1 Introduction
Since the discovery of superconductivity[1], people immediately realized that superconductors may be applied in many ways, thus enormous efforts have been dedicated to explore practical superconductors that meet the industrial requirements. Superconductors can be categorized into low-temperature superconductors (LTS) and high-temperature superconductors (HTS) according to their critical temperatures. Applicable low-temperature superconductors include NbTi, Nb3Sn and so on. NbTi superconductors have very good ductility and can be directly made into long wires[2]. Nb3Sn, on the other hand, due to its brittleness and hardness, is currently prepared by the bronze process and powder-in-tube (PIT) method[3]. Both superconductors are currently commercialized. The critical temperature of some cuprate high temperature superconductors is higher than the liquid nitrogen boiling temperature (77 K), such as Hg, Tl, Bi and RE-based systems[4, 5, 6, 7]. Due to the existence of toxic elements and the high anisotropy, the application of Hg-based and Tl-based systems is limited. Bi-based system, as a ceramic material, can not be directly made into wires and are mainly manufactured by powder-in-tube method, known as the first generation of high-temperature superconducting tapes. Because of its strong anisotropy, the vortex motion is very strong leading to a rather low irreversibility magnetic field in the liquid nitrogen temperature region, and the critical current density also decreases quickly in a small magnetic field at 77 K[8]. The critical temperature of REBa2Cu3O7-δ (REBCO) is about 90 K, and the anisotropy is relatively small compared with the Bi-system[9], which means that there is a high irreversibility field in the liquid nitrogen temperature region. However, due to the very small coherence length and the small tolerant misorientation angle between neighboring grains for the superconducting current to go through, the PIT technique was not successful to make the long wires for REBCO. In order to eliminate the weak links from the grain boundaries and obtain a better grain alignment to meet the requirements for practical applications, REBCO can be fabricated by epitaxial thin film method such as metal-organic chemical vapour deposition(MOCVD) and pulsed laser deposition (PLD)[10], etc. Thus the REBCO is known as the second generation of high-temperature superconducting tapes. However, the critical temperature of REBCO is about 90 K, lower than other systems with three or four layers of $CuO_{2}$ planes in the superconducting block. There is an another high-temperature superconducting system, (Cu,C)Ba2Can-1CunO2n+3 (n=1,2,3, …) [11, 12, 13, 14, 15]. Here (Cu,C)Ba2Ca2Cu3O9±δ((Cu,C)-1223) and (Cu,C)Ba2Ca3Cu4O11±δ((Cu,C)-1234) have a superconducting transition temperature of about 120 K and a higher irreversibility field than REBCO. But (Cu,C)-1223 and (Cu,C)-1234 bulks can only be synthesized by high-pressure methods, strongly limiting their applications. The growth of (Cu,C)-based superconducting films with high transition temperature is still challenging. The transition temperatures of (Cu,C)-1223 and (Cu,C)-1234 films grown on single-crystal substrates have reached 99.7 K and 96 K  [16, 17, 18], respectively. Similar to REBCO, these materials need to be epitaxially grown on flexible metallic substrate to meet the demands of industrial applications. However, to the best of our knowledge, up to now, there are no reports of (Cu,C)-1223 films grown on flexible metallic substrates. Due to its superior properties, it is important to grow this material on flexible metallic substrate in order to meet the demand for the industrial applications.
@@ 2 Experimental details
The (Cu,C)Ba2Ca2Cu3O9±δ thin films were epitaxially grown on Hastelloy tapes with LaMnO3 and CeO2 top buffer layer by pulsed laser deposition technique with a 248-nm KrF excimer laser, using the nominal composition target of Ba2Ca3Cu4.6Oy. The Hastelloy tapes with LaMnO3 and CeO2 top buffer layer are obtained from Companies of Suzhou Advanced Materials and Shanghai Superconductor, respectively. Fig. 1(a) and (b) show the architectures of the (Cu,C)-1223 coated conductors on different top layers. During deposition, a mixed gas of O2 and CO2 was introduced into the chamber, and the total pressure was maintained at 15-25 Pa. The deposition temperature for all films was controlled in the region 620-680 $\rm{{}^{\circ}C}$. The laser energy density of 1-2 J/cm2 and repetition rate of 5 Hz were used to ablate the target. After the deposition, the films were cooled down to room temperature in deposition pressure at a rate of $5\rm{{}^{\circ}C}/min$.
Figure 1: (a), (b) Sketch of two-type architectures of the (Cu,C)-1223 films. (c), (d) X-ray diffraction (XRD) patterns of the (Cu,C)Ba2Ca2Cu3O9±δ films on flexible Hastelloy substrates with the top layer of LaMnO3 and CeO2, respectively. (e), (f) AFM images of the (Cu,C)-1223/LaMnO3 film and the (Cu,C)-1223/CeO2 film.
@@ 3 Results and discussion
The x-ray diffraction patterns of the (Cu,C)-1223 films on flexible Hastelloy substrates with LaMnO3 and CeO2 top buffer layer are displayed in Fig. 1(c) and (d). All observed (00l) diffraction peaks can be indexed to (Cu,C)-1223 and the substrates, indicating that the films have been grown with c-axis orientation. Analysis of the XRD data yields a lattice constant $c$ = 14.82 Å for both films. Surface morphology and roughness have been investigated through atomic force microscopy (AFM) as shown in Fig. 1(e) and (f). The average roughness (Ra) of the (Cu,C)-1223/LaMnO3 film is about 1.9 nm, determined over the scan area $5\times 5\mu m^{2}$. For the (Cu,C)-1223/CeO2 film, the average roughness is about 5.0 nm. One possible reason for the obvious different $R_{a}$ is that the growth of the (Cu,C)-1223 films (a=3.86 Å) may be affected by epitaxial tensile strain from LaMnO3 (a=3.94 Å) and epitaxial compressive strain from CeO2 (a=3.82 Å).
Figure 2: (a), (b) Temperature dependent resistivity of the (Cu,C)Ba2Ca2Cu3O9±δ films on flexible metallic substrates with the top layer of LaMnO3 and CeO2. (c), (d) The temperature dependence of magnetic susceptibility under a magnetic field of 1 Oe.
@@
Figure 3: (a), (b) Temperature dependent resistivity of the (Cu,C)-1223/LaMnO3 film under magnetic fields $H\parallel ab$ and $H\parallel c$, respectively. (c) Temperature dependence of the upper critical field $H_{c2}$. The solid lines show the linear fitting results. (d) The resistivity as a function of the effective magnetic field at 90 K. The inset shows that angle dependent resistivity under different magnetic fields.